### v2

# Joint Probability Distribution of mRNA and Protein Molecules in a Stochastic Gene Expression Model

Abstract
Stochastic modeling of gene expression is a classic problem in theoretical biophysics. However, models formulated via chemical master equation have long been considered analytically intractable unless burst approximation is applied. This article shows that general stochastic gene expression models with an arbitrary number of gene states admit direct analysis. Based on chemical master equation and high-dimensional binomial moment method, we derive recurrence relations for binomial moments in steady state, yielding analytical expressions to arbitrary order in a hierarchical manner. Subsequently, the joint probability mass function of mRNA and protein copy number can be reconstructed. An algorithm is developed for numerical computation. Particularly, explicit expressions for low-order cumulants are presented. Compared with models under burst approximation, the mean remains exact, whereas the variance typically differs. We estimate the difference between two second-order binomial moments using functional analysis, therefore evaluating the validity of burst approximation.
1 Introduction
Gene expression is one of the most fundamental processes in molecular biology, where the genetic information is realized as functional gene products. In most cases, gene expression consists of transcription followed by translation. During transcription, messenger RNA (mRNA) is synthesized by copying specific segments of DNA, whereas translation is the process of producing proteins using mRNA as templates. To quantitatively understand and predict this process, constructing reasonable mathematical models is necessary. Notably, intrinsic stochasticity is the key feature of gene expression. Such inherent stochasticity originates from random collisions among molecules, and becomes relatively significant due to low numbers of mRNA and protein molecules in most cells [Centraldogmasinglemolecule2011Nature]. Randomized expression of genes can be beneficial for biological adaptation to changing environments [Noisegeneexpression2005Science]. However, how fluctuations propagate within and across cells, and are eventually brought under control, remains an open question.
In general, our primary concern is the mRNA and protein copy number during gene expression, whose dynamics should be modeled as a stochastic process. Note that deterministic approaches like reaction rate equation [AppliedStochasticAnalysis2019] certainly fail because of pronounced stochasticity. More specifically, this stochastic process should be a continuous-time process with a discrete state space. The state space should be discrete because the mRNA and protein counts are nonnegative integers. The simplest class of stochastic processes satisfying both conditions consists of continuous-time Markov chains, whose Kolmogorov forward equation is commonly termed chemical master equation [AppliedStochasticAnalysis2019, StochasticProcessesPhysics2007] in chemical physics literature. In general, chemical master equation is an infinite-dimensional coupled system of ordinary differential equations, and can be equivalently converted into a finite-dimensional partial differential equation system using generating function method or Laplace transform. Therefore, when analyzing a dynamical system described by chemical master equation, neither a universal treatment nor an effective general theory exists, and one must devise techniques tailored to each problem. In particular, compact analytical results are rarely anticipated, and in many cases different kinds of approximation techniques are applied [Approximationinferencemethods2017J.Phys.A-Math.Theor.]. The most widely-used approximation method is diffusion approximation [AppliedStochasticAnalysis2019], which approximates the chemical master equation with a Markov-type stochastic differential equation [BrownianMotionStochastic1998], conventionally termed chemical Langevin equation [chemicalLangevinequation2000J.Chem.Phys.]. This procedure approximates a continuous-time Markov chain with a diffusion process. Therefore, diffusion approximation also becomes inappropriate when the numbers of molecules in the system are extremely low, as the discreteness of the state space becomes significant. Despite the challenges, gene expression models based on chemical master equation are worth in-depth analysis given their biological importance.
Proposing stochastic gene expression models is straightforward. However, it has long been believed that complete gene expression models are analytically intractable. The simplest two-stage model is analyzed in the classic paper [Analyticaldistributionsstochastic2008Proc.Natl.Acad.Sci.U.S.A.]. In the two-stage model, mRNA molecules are continuously produced according to a Poisson process, and translation, hydrolysis of mRNA, and hydrolysis of protein all occur as memoryless, single-step reactions. Surprisingly, in this toy model, analytical distribution of protein copy number is infeasible, since the expression of the generating function contains an incomplete gamma function that cannot be reduced [Analyticaldistributionsstochastic2008Proc.Natl.Acad.Sci.U.S.A.]. Afterwards, very few studies focus on the complete gene expression models. Thereby, existing results about stochastic gene expression models are confined to low-order moments of several toy models [Modelsstochasticgene2005Phys.LifeRev.].
However, an effective approximation technique exists, namely, the burst approximation [Modelsstochasticgene2005Phys.LifeRev., Analyticaldistributionsstochastic2008Proc.Natl.Acad.Sci.U.S.A.]. The burst approximation builds upon the experimental conclusion that mRNAs decay substantially faster than proteins in most cells [Analyticaldistributionsstochastic2008Proc.Natl.Acad.Sci.U.S.A.]. Under burst approximation, the originally gene expression models can be greatly simplified and analytical results become available [Exactdistributionsstochastic2022Phys.Rev.E, Analyticaltimedependentdistributions2023SIAMJ.Appl.Math., Analyticaldistributionsstochastic2008Proc.Natl.Acad.Sci.U.S.A.]. Beyond their roles as approximate substitutes, stochastic gene expression models with translational bursting have their independent theoretical value. Particularly, models with multiple gene states under burst approximation have recently been studied using one-dimensional binomial moment method [Exactdistributionsstochastic2022Phys.Rev.E], and non-Markovian models can also be studied using techniques from queueing theory [Intrinsicnoisestochastic2011Phys.Rev.Lett.]. According to the equivalence between queueing system and gene expression model established in [Solvingstochasticgeneexpression2024Biophys.J.], models under burst approximation can be interpreted as queueing systems with batch arrivals. This perspective may explain the substantial analytical simplifications introduced by the burst approximation, and clarify the theoretical significance of related research. Nevertheless, whether the errors introduced by the burst approximation are controllable has not received much attention. The validity of burst approximation is only evaluated in the simplest two-stage model [Analyticaldistributionsstochastic2008Proc.Natl.Acad.Sci.U.S.A.], and validation for general models also motivates us to gain deeper analysis into complete gene expression models.
Additionally, general numerical methods exist for reaction systems described by chemical master equation. Stochastic simulation algorithm (SSA) [generalmethodnumerically1976J.Comput.Phys.] is a numerical method generating the sample paths of a given continuous-time Markov chain, one realization at a time. Multiple variants of SSA are also available [Stochasticsimulationchemical2007Annu.Rev.Phys.Chem.]. In this article, we employ Python package GillesPy2[GillesPy2biochemicalmodeling2023Lett.Biomath.] to perform SSA. Finite state projection algorithm (FSP) [finitestateprojection2006J.Chem.Phys.] is also a widely-used numerical method, which truncates the chemical master equation and reduces the problem to numerically solving a standard ordinary differential equation system. In general, numerical methods like SSA and FSP are computationally expensive and quickly become infeasible for large systems. Moreover, the aforementioned numerical methods cannot be used to derive further theoretical results. As a result, theoretically analyzing complete stochastic gene expression models is still important.
In this article, we demonstrate that gene expression models can be effectively analyzed without applying burst approximation. Using high-dimensional binomial moment method, we derive recurrence relations for binomial moments of mRNA and protein copy number, yielding exact analytical expressions for binomial moments of arbitrary order. We also design an algorithm that numerically computes, with high accuracy, all binomial moments in a hierarchical way. When binomial moments are obtained, joint probability mass function of mRNA and protein copy number can be constructed directly. In particular, we present the explicit expressions of low-order cumulants and binomial moments. Compared with models under burst approximation, the first-order binomial moment remains exact, whereas the second-order binomial moment generally differs. An upper bound is given for the difference between the second-order binomial moments obtained from two models, mainly using techniques from functional analysis. Additionally, we present the analytical expression of the multi-variable generating function by solving a partial differential equation system.
The article is structured as follows. In subsection 2.1, a general stochastic gene expression model is introduced, in which the gene has arbitrarily many states. In subsection 2.2, the chemical master equation describing the dynamics is presented, and then converted to a partial differential equation system using standard generating function method. Using high-dimensional binomial moment method, an ordinary differential equation hierarchy governing the evolution of binomial moments of mRNA and protein copy number is derived. We present the concise recurrence relations in steady state and design an algorithm for numerical computation. These results are presented in subsection 3.1. In subsection 3.2, we provide an equality to reconstruct the probability mass function from binomial moments. Analytical expressions for low-order cumulants of the number of mRNA and protein molecules are given in subsection 3.3, and are compared with those from models under burst approximation in section 4.
2 Stochastic Gene Expression Model
2.1 Model Description
We now introduce a general stochastic model of gene expression, formulated as a chemical reaction system schemed in (1). In this model, the gene we consider is assumed to have $N$ different states, namely, $\mathcal{S}_{i}\;(1\leq i\leq N)$, and the gene transits arbitrarily among these states in a Markovian manner. In each state, mRNA molecules are transcribed in a different rate and gene-state is allowed to switch upon transcription. Transcribed mRNA molecules are confronted with competing reaction pathways of hydrolysis and translation. Hydrolysis of mRNA molecules occurs at constant rate, independent of the state of gene or number of protein molecules. Specific type of protein molecules are translated from living mRNA molecules at constant rate, and also undergo hydrolysis once produced. Based on the general theory of stochastic chemical reaction kinetics [AppliedStochasticAnalysis2019], (1) defines a continuous-time Markov chain $(S(t),M_{1}(t),M_{2}(t))_{t\geq 0}$ with state space $\{1,2,\cdots,N\}\times\mathbb{N}\times\mathbb{N}$, where $S(t)$, $M_{1}(t)$ and $M_{2}(t)$ denote the state of gene, the number of mRNA molecules in the system, and the number of protein molecules in the system.
[…]
$\rightarrowfill@\displaystyle{}\mathrm{{Protein}}{}\mathrel{\mathop{\makebox[0.0pt]{\rightarrowfill@\displaystyle}}\limits^{\mkern 5.0mu{}\mathrm{{}}\mkern 9.0mu}}{}{}\mathrm{\text{$\emptyset$}}$
Denote by $a_{i,j}\;(i\neq j,\;1\leq i,j\leq N)$ the transition rates among the states of gene without transcription, and $b_{i,j}\;(1\leq i,j\leq N)$ the transition rates with production of one mRNA molecule. Let $v$, $u$, and $\delta$ denote the translation rate, the mRNA degradation rate, and the protein degradation rate, respectively. Define $a_{i,i}:=-\sum_{\begin{subarray}{c}k=1\\ k\neq i\end{subarray}}^{N}a_{i,k}-\sum_{k=1}^{N}b_{i,k}$, $D_{0}:=(a_{i,j})_{N\times N}$, $D_{1}:=(b_{i,j})_{N\times N}$, and $D:=D_{0}+D_{1}$.
Detailed stochastic gene expression models tracking the dynamics of both mRNA and protein counts are rarely studied before, because they have long been assumed analytically intractable even in the simplest case. The simplest special case in (1), namely the case where $N=1$, is first analyzed in [Analyticaldistributionsstochastic2008Proc.Natl.Acad.Sci.U.S.A.]. Using generating function method and the method of characteristics for solving linear partial differential equations, the authors of [Analyticaldistributionsstochastic2008Proc.Natl.Acad.Sci.U.S.A.] obtain analytical expression of the generating function but find that it involves an integral that cannot be calculated, namely, $\int\frac{\mathrm{e}^{x}}{x}$. Therefore, they stop further analyzing this model and propose an approximation technique, the burst approximation, that is now predominantly used by the community. As a result, the detailed stochastic gene expression model (1) has long been assumed intractable and has never been thoroughly studied even in any simple cases.
2.2 Chemical Master Equation
Assume there are no mRNA or protein molecules in the system at time 0 and the gene is in a given state, say, $S_{i}\;(1\leq i\leq N)$. Denote by $\mathbb{P}_{i,j}(m,n;t)\;(1\leq j\leq N,m\in\mathbb{N},n\in\mathbb{N},t\geq 0)$ the probability that at time $t$ there are $m$ mRNA molecules, $n$ protein molecules and the gene is in state $S_{j}$. Define $\mathbb{P}(m,n;t)$ to be a $N\times N$ matrix whose $(i,j)$-th element is $\mathbb{P}_{i,j}(m,n;t)$. The chemical master equation of (1) is
$\frac{\partial}{\partial t}\mathbb{P}_{i,j}(m,n;t)$ $=\sum_{\begin{subarray}{c}s=1\\ s\neq j\end{subarray}}^{N}a_{s,j}\mathbb{P}_{i,s}(m,n;t)+\sum_{s=1}^{N}b_{s,j}\mathbb{P}_{i,s}(m-1,n;t)$ (2)
$+mv\mathbb{P}_{i,j}(m,n-1;t)+(m+1)u\mathbb{P}_{i,j}(m+1,n;t)$
$+(n+1)\delta\mathbb{P}_{i,j}(m,n+1;t)-(mu+n\delta+mv)\mathbb{P}_{i,j}(m,n;t)$
$-(\sum_{\begin{subarray}{c}s=1\\ s\neq j\end{subarray}}^{N}a_{j,s}+\sum_{s=1}^{N}b_{j,s})\mathbb{P}_{i,j}(m,n;t).$
Using standard generating function method, or equivalently, Laplace transform, we define the matrix-form generating function of $\mathbb{P}(m,n,t)$ as
$\mathcal{G}(z,w;t)=\sum_{m=0}^{\infty}\sum_{n=0}^{\infty}z^{m}w^{n}\mathbb{P}(m,n;t),\;\;z,w\in\mathbb{R},\;\lvert z\rvert\leq 1,\;\lvert w\rvert\leq 1.$ (3)
The infinite-dimensional ordinary differential equation system (2) can then be transformed into a finite-dimensional partial differential equation system, namely,
$\frac{\partial}{\partial t}\mathcal{G}(z,w;t)=$ $\mathcal{G}(z,w;t)D_{0}+z\mathcal{G}(z,w;t)D_{1}$ (4)
$+\left[u(1-z)+vz(w-1)\right]\frac{\partial}{\partial z}\mathcal{G}(z,w;t)+\delta(1-w)\frac{\partial}{\partial w}\mathcal{G}(z,w;t).$
Detailed derivation can be found in the Supplementary Material.
Remarkably, by setting $w=1$, we focus on the marginal distribution of mRNA copy number, and (4) reduces to
$\frac{\partial}{\partial t}\mathcal{G}(z,1;t)=$ $\mathcal{G}(z,1;t)D_{0}+z\mathcal{G}(z,1;t)D_{1}+u(1-z)\frac{\partial}{\partial z}\mathcal{G}(z,1;t),$ (5)
which is exactly the partial differential equation system appeared in our previous work [lu2025stochastickineticsmrnamolecules]. This can be intuitively understood since the fluctuations of protein copy number does not contribute to the fluctuation at mRNA level, while changes in mRNA copy number do affect fluctuations at protein level. This phenomena is widely-known as dynamic disorder [Singlemoleculeapproachdispersed2002J.Chem.Phys.]. Additionally, the probability distribution of protein copy number cannot be treated marginally; instead, the joint probability distribution needs to be considered.
Actually, we also construct the analytical solution to the partial differential equation system (5) using the method of characteristics, which can be found in the Supplementary Material. We note that further analysis based on the analytical solution requires additional development and is not pursued in this article.
3 Reconstruct Probability Mass Function from Binomial Moments
3.1 Two-dimensional Binomial Moment Method
In this and the following section, we aim to establish an efficient numerical scheme for computing the steady-state joint distribution of mRNA and protein copy number. Our approach builds on standard binomial moment method [Binomialmomentequations2011Phys.Rev.Lett., Stochasticanalysiscomplex2012Phys.Rev.Ea, momentconvergencemethodstochastic2016J.Chem.Phys.a], which analyzes the moments of a probability distribution and subsequently reconstructs the corresponding probability mass function. We first obtain the binomial moments in this section.
Define the binomial moment of the joint probability distribution as
$\mathcal{B}_{p,q}(t)=\sum_{m=p}^{\infty}\sum_{n=q}^{\infty}\binom{m}{p}\binom{n}{q}\mathbb{P}(m,n;t),\;p\in\mathbb{N},\;q\in\mathbb{N},$ (6)
where $\binom{m}{p}$ denotes the combinatorial coefficient.
Binomial moments are linear combinations of moments defined by $\langle M_{1}(t)^{p}M_{2}(t)^{q}\rangle$
$\sum_{m=0}^{\infty}\sum_{n=0}^{\infty}m^{p}n^{q}\mathbb{P}(m,n;t)$ [CourseProbabilityTheory2000], therefore binomial moments exist if and only if moments of arbitrary orders exist for each element in $\mathbb{P}(m,n;t)$. The name “binomial” comes from combinatorial coefficients in (6). Additionally, denote the stationary binomial moments by $\mathcal{B}_{p,q}:=\lim_{t\rightarrow\infty}\mathcal{B}_{p,q}(t)$ and the stationary probability mass function by $\mathbb{P}(m,n):=\lim_{t\rightarrow\infty}\mathbb{P}(m,n;t)$.
According to the definition (6), we have
$\mathcal{G}(z,w;t)$ $\equiv\sum_{m=0}^{\infty}\sum_{n=0}^{\infty}z^{m}w^{n}\mathbb{P}(m,n;t)=\sum_{p=0}^{\infty}\sum_{q=0}^{\infty}s^{p}r^{q}\mathcal{B}_{p,q}(t),$ (7)
where s:=z1s:=z-1 and r:=w1r:=w-1. Substituting (7) into (4), we may obtain an ordinary differential equation hierarchy for $\mathcal{B}_{p,q}(t)$. Setting the time-derivatives to zero, we get a linear system for binomial moments at steady state. Specifically, we have
$\mathcal{B}_{p,q}\left(D-up\bm{I}_{N}-\delta q\bm{I}_{N}\right)+v(p+1)\mathcal{B}_{p+1,q-1}+vp\mathcal{B}_{p,q-1}+\mathcal{B}_{p-1,q}D_{1}=0,\;p\in\mathbb{N},\;q\in\mathbb{N}.$ (8)
Note that $\mathcal{B}_{0,0}=\bm{I}_{N}$ by definition and $\mathcal{B}_{p,q}$ is taken as $\bm{0}_{N\times N}$ if $p<0$ or $q<0$. See Supplementary Material for the detailed derivation.
Arbitrary-order binomial moments can be obtained in a hierarchical manner using (8). Define the layer, or order, of a binomial moment $\mathcal{B}_{p,q}$ as $L:=p+q$. All binomial moments are assembled according to their layers, proceeding from lower to higher layers. Within each layer, binomial moments are ordered by increasing $p$ and derived sequentially from right to left. This overall procedure is illustrated in Figure 1, and the detailed algorithmic process is presented in Algorithm S.1.
Figure 1: Hierarchy of Binomial Moments: This image illustrates the hierarchical structure for calculating binomial moments up to desired orders. The layers are organized from $L=0$ at the top to $L=5$ at the bottom. Within each layer, binomial moments are arranged from left to right in order of increasing $p$ (and correspondingly decreasing $q$). The calculation proceeds from lower to higher layers. Within each layer, binomial moments are derived sequentially from right to left. The arrows indicate the dependency flow, showing how binomial moments in higher layers depend on values calculated earlier in the same or preceding layers.
Recall that $D$ is the generator of the Markov chain $(S(t))_{t\geq 0}$. Thereby, $\gamma\bm{I}_{N}-D$ is a strictly diagonally dominant matrix for any $\gamma>0$, and is nonsingular by Lévy-Desplanques theorem [MatrixAnalysis2012]. Additionally, strict diagonal dominance guarantees the numerical stability while performing a LU factorization [MatrixComputations2013]. Thus, the numerical scheme Algorithm S.1 yields binomial moments with high accuracy.
Note that analytical expressions of the binomial moments of mRNA copy number, namely, $\mathcal{B}_{p,0}\;(p\in\mathbb{N})$, agrees with existing results [Multimodalityflexibilitystochastic2013Bull.Math.Biol., lu2025stochastickineticsmrnamolecules, Promotermediatedtranscriptionaldynamics2014Biophys.J.]. However, the derivation here is primarily formal in comparison with our earlier work [lu2025stochastickineticsmrnamolecules], in which the time-dependent binomial moments are derived based on the analytical solution to the chemical master equation, and the stationary expressions are obtained by taking temporal limit.
3.2 Reconstruct Probability Mass Function
Under certain condition, all moments can uniquely determine the underlying probability distribution [CourseProbabilityTheory2000]. Particularly, the binomial moments may determine the probability mass function through a compact identity. According to (7), stationary probability mass function $\mathbb{P}(m,n)$ can be reconstructed from binomial moments according to
$\mathbb{P}(m,n)=\sum_{p=m}^{\infty}\sum_{q=n}^{\infty}(-1)^{p+q+m+n}\binom{p}{m}\binom{q}{n}\mathcal{B}_{p,q}.$ (9)
See Supplementary Material for the detailed derivation.
We note that, in practice, computation based on (9) suffers from severe floating-point error if not implemented properly. This is because for large $p$ and $q$, multiplying the relatively small binomial moment with the extremely large combinatorial coefficients can introduce significant floating-point errors. This is essentially the same numerical issue encountered and analyzed in [Exactlysolvablemodels2020J.Chem.Phys., lu2025stochastickineticsmrnamolecules]. In the accompanying codes, we take the approach in [lu2025stochastickineticsmrnamolecules], carrying out computations in the logarithmic domain and exponentiating the final results.
We analyze an example using both the analytical results (8), SSA, and FSP; the results are shown in Figure 2, Figure 3, and Figure 4, respectively. The mean squared error between the joint probability mass function in Figure 2 and Figure 3 is $1.9208\times 10^{-8}$; and between the joint probability mass function in Figure 2 and Figure 4 is $1.7685\times 10^{-12}$. See the accompanying Jupyter Notebook for detail.
Figure 2: Probability Distribution of mRNA and Protein Copy Number obtained using Analytical Results: This stem plot illustrates the probability mass function obtained by first computing binomial moments according to Algorithm S.1, and then reconstructing via (9). In this example, parameters in the model (2) are $D_{0}=$ $\begin{pmatrix}-2.02&0.01&0.01\\ 0.1&-7.2&0.1\\ 0&0.01&-6.01\end{pmatrix}$ , $D_{1}=$ $\begin{pmatrix}1&0&1\\ 1&5&1\\ 0&1&5\end{pmatrix}$ , $u=3$, $v=2$, and $\delta=1$. The largest layer in Algorithm S.1 is $L_{\text{max}}=300$. $\mathbb{P}(m,n)$ is computed for $0\leq m,n\leq 16$. The infinite series (9) is truncated up to the largest layer $L_{\text{max}}$. In the stem plot in the top panel, we use the coarse-grained probability mass function $P(m,n)$ defined in subsection $3.$ The two stem plots in the bottom panel are the marginal distribution of mRNA and protein copy number, respectively. The marginal distributions are obtained directly from the joint probability distribution.
Figure 3: Probability Distribution of mRNA and Protein Copy Number obtained through Stochastic Simulations: The parameters in the model (2) are the same as those in Figure 2. The histogram in the upper panel is generated with $1\times 10^{6}$ trajectories of SSA, all truncated at dimensionless time $t=50$. The initial condition is $S(0)=1$, $M_{1}(0)=0$, and $M_{2}(0)=0$. Python package GillesPy2 is implemented with C++ solver. The two stem plots in the bottom panel are the marginal distribution of mRNA and protein copy number, respectively. The marginal distributions are obtained directly from the joint probability distribution.
Figure 4: Probability Distribution of mRNA and Protein Copy Number obtained through Finite State Projection Algorithm: The parameters in the model (2) are the same as those in Figure 2. The histograms in the top panel (joint probability distribution of mRNA and protein copy number) are plotted according to FSP at $t=20$, and the truncation error is below $1\times 10^{-4}$. The initial condition is $S(0)=1$, $M_{1}(0)=0$, and $M_{2}(0)=0$. The two stem plots in the bottom panel are the marginal distribution of mRNA and protein copy number, respectively. The marginal distributions are obtained directly from the joint probability distribution.
3.3 Low-Order Moments
In this section, we present analytical expressions for several important low-order binomial moments, and first-order and second-order cumulants of the joint probability distribution of mRNA and protein copy number.
Using the hierarchal approach of calculating the binomial moments described in the above section, we can readily obtain explicit expressions of low-order binomial moments. For simplicity, we now stop tracking the state of gene by considering only the coarse-grained binomial moments $B_{p,q}:=\bm{\pi}^{\top}\mathcal{B}_{p,q}\bm{1}$ and the probability mass function $P(m,n):=\bm{\pi}^{\top}\mathbb{P}(m,n)\bm{1}$ for $p,q,m,n\in\mathbb{N}$, similar to [lu2025stochastickineticsmrnamolecules]. $\bm{1}\in\mathbb{R}^{N\times 1}$ denotes the all-ones vector. $\bm{\pi}\in\mathbb{R}^{N\times 1}$ is the invariant distribution of the underlying Markov chain $(S(t))_{t\geq 0}$ characterized by $D$. Assume $D$ is irreducible. Then $\bm{\pi}$ is the unique vector satisfying $\bm{\pi}^{\top}D=\bm{0}_{1\times N}$ and $\bm{\pi}^{\top}\bm{1}=1$. Explicit expressions of $\mathcal{B}_{p,q}$, $B_{p,q}$ with $L\leq 2$ and the detailed derivation can be found in the Supplementary Material. For comparison with binomial moments derived from models under burst approximation in section 4, we particularly present here the first two binomial moments of protein copy number.
$B_{0,1}=\frac{v}{u\delta}\bm{\pi}^{\top}D_{1}\bm{1},$ (10)
$B_{0,2}=\frac{v^{2}}{2u\delta(u+\delta)}\bm{\pi}^{\top}D_{1}\bm{1}+\frac{v^{2}}{2\delta(u+\delta)}\bm{\pi}^{\top}D_{1}(u\bm{I}_{N}-D)^{-1}(\delta\bm{I}_{N}-D)^{-1}D_{1}\bm{1}$
$\;\;\;\;\;\;\;\;\;\;\;\;\;+\frac{v^{2}}{2u\delta(u+\delta)}\bm{\pi}^{\top}D_{1}(u\bm{I}_{N}-D)^{-1}D_{1}\bm{1}.$
Based on explicit expressions of low-order binomial moments, we can further calculate several key statistical quantities of the joint probability distribution of mRNA and protein count at steady state. Specifically, we calculate the first-order and second-order cumulants in the Supplementary Material.
4 Accuracy of Burst Approximation
In this section, we evaluate the validity of the well-known burst approximation by comparing the analytical expressions of low-order binomial moments obtained before and after applying the burst approximation.
Burst approximation is commonly used to simplify the complete gene expression model (1). Intuitively, burst approximation aims to capture the experimental phenomena that proteins are often produced in bursts. In one burst, proteins are synthesized at a relatively high rate over a short interval, followed by a long period of silence. One important quantitative relation underlying such burst phenomena is that, in general, mRNAs have much shorter lifetimes than proteins. In the notations of (1), this corresponds to $u\gg\delta$. However, it has been argued that $u\gg\delta$ is far from sufficient to generate genuine translation bursts, and bust approximation is better understood as a conceptual mathematical technique [Modelsstochasticgene2005Phys.LifeRev.].
Under burst approximation, mRNA molecules can be eliminated formally from the system, and only the one-dimensional probability distribution of protein molecules needs to be revealed. As a result, stochastic models under burst approximation are typically easier to analyze. For models with one active gene state and multiple inactive states, a compact analytical expression for the probability mass function can be derived in terms of the generalized hypergeometric function [Exactdistributionsstochastic2022Phys.Rev.E, Analyticaldistributionsstochastic2008Proc.Natl.Acad.Sci.U.S.A.]. By contrast, this is generally impossible for the complete gene expression model (2) according to the analytical expression of the generating function (See the Supplementary Material). General models with multiple gene states are also studied under burst approximation, and a recurrence relation for binomial moments can also be obtained using one-dimensional binomial moment method [Exactdistributionsstochastic2022Phys.Rev.E]. Additionally, non-Markovian models (models that are not continuous-time Markov chains) under burst approximation exist and can be studied using queueing theory [Intrinsicnoisestochastic2011Phys.Rev.Lett.]. We note that, according to the equivalence between queueing system and gene expression model established in [Solvingstochasticgeneexpression2024Biophys.J.], models under burst approximation can be interpreted as queueing systems with batch arrivals [Analysisinfiniteserverqueue2002QueueingSyst.]. This perspective may explain the significant convenience introduced by the burst approximation.
In general, stochastic gene expression models under burst approximation possess independent theoretical value. However, in this article, we focus on the validity of burst approximation when replacing the original reaction system (1) with an alternative. Therefore, the general setting is not introduced here. A systematic analysis of general stochastic gene expression models under burst approximation will be reported in a forthcoming paper [lu2025stochastickineticsproteinmolecules].
Here, we approximate (1) with the following reaction system (11).
[…]
$\rightarrowfill@\displaystyle{}\mathrm{{Protein}}{}\mathrel{\mathop{\makebox[0.0pt]{\rightarrowfill@\displaystyle}}\limits^{\mkern 5.0mu{}\mathrm{{}}\mkern 9.0mu}}{}{}\mathrm{\text{$\emptyset$}}$
In (11), the species ${S}_{i}\;(1\leq i\leq N)$, Protein, and the parameters $a_{i,j}\;(i\neq j,\;\;1\leq i,j\leq N)$, $\delta$ carry the same meaning as in (1). Additionally, under the notations in (1), $b_{i,j}^{[r]}\;(1\leq i,j\leq N,\;\;r=1,2,\cdots)$ are given by the following relation.
$b_{i,j}^{[r]}=\left(\frac{v}{u+v}\right)^{r}\left(1-\frac{v}{u+v}\right)b_{i,j},1\leq i,j\leq N,\;\;r=1,2,\cdots.$ (12)
The relation (12) can be interpreted as follows. Once transcribed, an mRNA molecule is subject to two competing reaction pathways, namely, hydrolysis and translation. More specifically, this can be seen as the competing binding of decay complexes that promote hydrolysis, and recruitment of initiation factors that engage the ribosome for translation. Since the probability of initiating translation rather than hydrolysis is $v/(u+v)$, the number of protein molecules produced from a single mRNA molecule follows a geometric distribution with parameter $u/(u+v)$. Hence, (12) readily follows. We note that the geometrically distributed burst size is consistent with experimental observations [Stochasticproteinexpression2006Naturea].
With the notations introduced before, the first-order and second-order binomial moments of protein copy number in (11) are
$\widetilde{B_{1}}$ $=\frac{v}{u\delta}\bm{\pi}^{\top}D_{1}\bm{1},$ (13)
$\widetilde{B_{2}}$ $=\frac{v^{2}}{2u^{2}\delta}\bm{\pi}^{\top}D_{1}\bm{1}+\frac{v^{2}}{2u^{2}\delta}\bm{\pi}^{\top}D_{1}\left(\delta\bm{I}_{N}-D\right)^{-1}D_{1}\bm{1}.$
The detailed derivation can be found in [lu2025stochastickineticsproteinmolecules].
Compared with (10), it is worth noting that the first-order binomial moment, equivalently the expectation of the protein copy number, remains exact under burst approximation. By contrast, the second-order binomial moment is explicitly altered when burst approximation is applied. In particular, an upper bound can be derived for the difference between the binomial moments obtained from the two models, using techniques from functional analysis and an analogue to Theorem 4.1.2 in [MatrixComputations2013]. Specifically, we have
$\lvert B_{0,2}-\widetilde{B_{2}}\rvert\leq$ $\frac{v^{2}}{2u^{2}(u+\delta)}\lVert D_{1}\rVert_{\infty}+\frac{v^{2}}{u^{2}\delta(u+\delta)}\lVert D_{1}\rVert_{\infty}^{2}+\frac{v^{2}}{2u^{2}\delta^{2}}\lVert D_{1}\rVert_{\infty}^{2}\lVert D\rVert_{\infty}.$ (14)
Note that the infinity norm of a matrix, denoted by $\lVert\cdot\rVert_{\infty}$, is the maximum absolute row sum of this matrix. The proof can be found in the Supplementary Material. According to (14), the difference converges to zero at the rate of $1/u^{2}$ if $u\rightarrow\infty$ while the other parameters are fixed. In general, $u/\delta\gg 1$ does not guarantee the validity of the buret approximation. In the left panel of Figure 5, both the upper bound given by (14) and the gap between variances converges to zero as $u$ grows and all other parameters remain the same. However, according to the right panel of Figure 5, although $u/\delta\gg 1$, the gap between variances of the protein copy number in complete gene expression models and the corresponding surrogate models can be arbitrarily large, indicating that the burst approximation can be inaccurate in some cases.
Figure 5: Gap between Variance and its Upper Bound: In the left panel, all the parameters expect for $u$ are the same as those in Figure 2 and are fixed. While $u$ varies in $[1,20]$, we compute the discrepancy between the variances of the protein copy number according to (10) and Equation 13, and also evaluate its upper bound given by (14). In the right panel, $u=10$, $v=2$, and $\delta=1$ are fixed, while $D_{0}$ and $D_{1}$ take the form $a_{i,j}=0\;(i\neq j)$ and $b_{i,j}=i$. For the order of the model $1\leq N\leq 100$, we compute the discrepancy between the variances of the protein copy number according to (10) and Equation 13, and also evaluate its upper bound given by (14). Note that both illustrations are plotted in $\ln(\cdot)$ scale, and that the discrepancy between variances equals two times the discrepancy between second-order binomial moments.
5 Discussion and Conclusion
In this paper, we establish an effective approach to analyzing stochastic gene expression models without resorting to burst approximation. Analytical expressions of binomial moments of mRNA and protein counts can be obtained up to any order at steady state in a hierarchical manner. Numerical computation of binomial moments is both fast and accurate. Subsequently, joint probability mass function can be reconstructed. In particular, based on analytical expressions of low-order cumulants, we can rigorously evaluate the validity of the burst approximation for general gene expression models.
We first note that asymptotic behavior of binomial moments and probability mass function needs further analysis. In [lu2025stochastickineticsmrnamolecules], we derive elegant upper bounds for binomial moments and probability mass function of mRNA copy number, which are $B_{p,0},\;p\in\mathbb{N}$ and $\sum_{n=0}^{\infty}P(m,n),\;m\in\mathbb{N}$ in this article, respectively. However, deriving concise upper bounds for $B_{p,q},\;p,q\in\mathbb{N}$ or $P(m,n),\;m,n\in\mathbb{N}$ appears substantially more involved. Note that asymptotic analysis is crucial to proving the convergence of (9) and designing an appropriate truncation strategy. A rigorous treatment is left to future work.
In the end, we examine the gene-state topology from a multiscale modeling perspective. Given the flexibility of our general model (1), it becomes essential to predefine the number of states $N$ and the transition structure, namely $D_{0}$ and $D_{1}$. Otherwise, the model may yield highly pathological results and nonphysical predictions. Fortunately, from a multiscale modeling perspective, this task may be accomplished by molecular dynamics (MD) simulations [StatisticalMechanicsTheory2023]. At the atomistic level, dynamics of macromolecules, such as DNA and proteins, can be accurately captured using MD simulations. Physically, each gene state corresponds to a metastable configuration of the underlying macromolecule, namely DNA in this problem. Thus, the Markov chain of gene state is a coarse-grained representation of the underlying dynamics. Designing the gene-state topology essentially amounts to constructing a Markov state model (MSM) from MD trajectories, which is a standard method for analyzing MD data and understanding conformational transitions [Identificationslowmolecular2013J.Chem.Phys., MetastabilityMarkovState2013]. Constructing a MSM typically involves two steps. The first step is to project the high-dimensional MD data, namely the coordinates of all atoms in a macromolecule, onto a relatively lower-dimensional latent space. The above dimensionality reduction relies on specifying collective variables (CVs), which map the complete configurations to low-dimensional representations but preserve non-trivial dynamical behaviors of the molecule. Next, one clusters the data in CV space and uses cluster centroids as discrete states of a Markov chain, whose transition rates are estimated at a given lag time. Nevertheless, constructing CVs remains a challenge for real-world macromolecules, although machine learning techniques have been incorporated to develop many useful methods [Collectivevariablediscovery2022RSCAdv., unifiedframeworkmachine2023J.Chem.Phys.]. Therefore, further work is needed to develop a physically grounded, bottom-up stochastic gene expression model.
Acknowledgments
Y.L. thanks Pinchen Xie and Ruiqi Gao for fruitful discussions. Y.L. was supported by Natural Science Foundation of China (125B10002), FDUROP (Fudan Undergraduate Research Opportunities Program) (24260), and Shanghai Undergraduate Training Program on Innovation and Entrepreneurship grant (S202510246532). Y.Z. was supported by National Key R&\&D Program of China (2024YFA1012401), the Science and Technology Commission of Shanghai Municipality (23JC1400501), and Natural Science Foundation of China (12241103).
Conflict of Interest
The authors have no conflicts to disclose.
Code Availability
The Python code will be released in a public repository soon.
Appendix A Matrix-form Chemical Master Equation and Generating Function Method
In this section, we show that the chemical master equation (2) can be equivalently reformulated as the partial differential equation system (4) using standard generating function method.
Recall the definition of $a_{i,i}$, the chemical master equation of (1), namely, (2), can be rearranged as
$\frac{\partial}{\partial t}\mathbb{P}_{i,j}(m,n;t)$ $=\sum_{s=1}^{N}a_{s,j}\mathbb{P}_{i,s}(m,n;t)+\sum_{s=1}^{N}b_{s,j}\mathbb{P}_{i,s}(m-1,n;t)$ (15)
$+mv\mathbb{P}_{i,j}(m,n-1;t)+(m+1)u\mathbb{P}_{i,j}(m+1,n;t)$
$+(n+1)\delta\mathbb{P}_{i,j}(m,n+1;t)-(mu+n\delta+mv)\mathbb{P}_{i,j}(m,n;t).$
Given $m$, $n$ and $t$, assemble the entries $\mathbb{P}_{i,j}(m,n;t)$ into a matrix according to the subindex $(i,j)$, obtaining the matrix-form probability mass function $\mathbb{P}(m,n;t)\in\mathbb{R}^{N\times N}$. With this notation, (15) takes the compact form
$\frac{\partial}{\partial t}\mathbb{P}(m,n;t)$ $=\mathbb{P}(m,n;t)D_{0}+\mathbb{P}(m-1,n;t)D_{1}$ (16)
$+mv\mathbb{P}(m,n-1;t)+(m+1)u\mathbb{P}(m+1,n;t)$
$+(n+1)\delta\mathbb{P}(m,n+1;t)-(mu+n\delta+mv)\mathbb{P}(m,n;t).$
Then we refer to standard generating function method. Recall that the matrix-form generating function $\mathcal{G}(z,w;t)$ is defined in the main text by (3). The chemical master equation (16) can be equivalently converted into the partial differential equation system.
Converting (15) into matrix-form and using generating function, the original chemical master equation system can be neatly expressed as the following partial differential equation system (4).
Appendix B Differential Equations for the Binomial Moments
In this section, we establish the relations between binomial moments and the probability mass function, and derive the hierarchy of ordinary differential equations governing binomial moments.
According to the formal definition (6) of binomial moments $\mathcal{B}_{p,q}(t)$, we can now verify the equality (7), stating that binomial moments are Taylor coefficients of the generating function $\mathcal{G}(z,w;t)$ expanded around $z=1,w=1$. Note that, since $z=s+1$ and $w=r+1$,
$\mathcal{G}(z,w;t)$ $\equiv\sum_{m=0}^{\infty}\sum_{n=0}^{\infty}z^{m}w^{n}\mathbb{P}(m,n;t)=\sum_{m=0}^{\infty}\sum_{n=0}^{\infty}(s+1)^{m}(r+1)^{n}\mathbb{P}(m,n;t)$ (17)
$=\sum_{m=0}^{\infty}\sum_{p=0}^{m}\sum_{n=0}^{\infty}\sum_{q=0}^{n}\binom{m}{p}\binom{n}{q}s^{p}r^{q}\mathbb{P}(m,n;t)$
$=\sum_{p=0}^{\infty}\sum_{m=p}^{\infty}\sum_{q=0}^{\infty}\sum_{n=q}^{\infty}\binom{m}{p}\binom{n}{q}s^{p}r^{q}\mathbb{P}(m,n;t)$
$=\sum_{p=0}^{\infty}\sum_{q=0}^{\infty}s^{p}r^{q}\left[\sum_{m=p}^{\infty}\sum_{n=q}^{\infty}\binom{m}{p}\binom{n}{q}\mathbb{P}(m,n;t)\right]$
$\equiv\sum_{p=0}^{\infty}\sum_{q=0}^{\infty}s^{p}r^{q}\mathcal{B}_{p,q}(t).$
Likewise, by inverting the above derivation, we arrive at the relation (9) showing how binomial moments determine the probability mass function.
$\sum_{p=0}^{\infty}\sum_{q=0}^{\infty}s^{p}r^{q}\mathcal{B}_{p,q}(t)$ $\equiv\sum_{p=0}^{\infty}\sum_{q=0}^{\infty}(z-1)^{p}(w-1)^{q}\mathcal{B}_{p,q}(t)$ (18)
$=\sum_{p=0}^{\infty}\sum_{m=0}^{p}\sum_{q=0}^{\infty}\sum_{n=0}^{q}(-1)^{p+q-m-n}\binom{p}{m}\binom{q}{n}z^{m}w^{n}\mathcal{B}_{p,q}(t)$
$=\sum_{m=0}^{\infty}\sum_{p=m}^{\infty}\sum_{n=0}^{\infty}\sum_{q=n}^{\infty}(-1)^{p+q-m-n}\binom{p}{m}\binom{q}{n}z^{m}w^{n}\mathcal{B}_{p,q}(t)$
$=\sum_{m=0}^{\infty}\sum_{n=0}^{\infty}z^{m}w^{n}\left[\sum_{p=m}^{\infty}\sum_{q=n}^{\infty}(-1)^{p+q-m-n}\binom{p}{m}\binom{q}{n}\mathcal{B}_{p,q}(t)\right]$
$=\sum_{m=0}^{\infty}\sum_{n=0}^{\infty}z^{m}w^{n}\left[\sum_{p=m}^{\infty}\sum_{q=n}^{\infty}(-1)^{p+q+m+n}\binom{p}{m}\binom{q}{n}\mathcal{B}_{p,q}(t)\right].$
By the identity given in (7), the above expression can equivalently be expressed as $\mathcal{G}(z,w;t)\equiv\sum_{m=0}^{\infty}\sum_{n=0}^{\infty}z^{m}w^{n}\mathbb{P}(m,n;t)$. From the uniqueness of the Taylor coefficients, it follows that
$\mathbb{P}(m,n;t)=\sum_{p=m}^{\infty}\sum_{q=n}^{\infty}(-1)^{p+q+m+n}\binom{p}{m}\binom{q}{n}\mathcal{B}_{p,q}(t).$ (19)
To obtain the differential equation system governing binomial moments, we now substitute (7) into the partial differential equation system (4), and it follows that
$\sum_{p=0}^{\infty}\sum_{q=0}^{\infty}s^{p}r^{q}$ $\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{B}_{p,q}(t)=\sum_{p=0}^{\infty}\sum_{q=0}^{\infty}s^{p}r^{q}\mathcal{B}_{p,q}(t)D_{0}+(s+1)\sum_{p=0}^{\infty}\sum_{q=0}^{\infty}s^{p}r^{q}\mathcal{B}_{p,q}(t)D_{1}$ (20)
$+\left[-us+v(s+1)r\right]p\sum_{p=1}^{\infty}\sum_{q=0}^{\infty}s^{p-1}r^{q}\mathcal{B}_{p,q}(t)-\delta rq\sum_{p=0}^{\infty}\sum_{q=1}^{\infty}s^{p}r^{q-1}\mathcal{B}_{p,q}(t).$
Arranging the right hand side of (20) according to the order of $s$ and $r$, we obtain a hierarchy of ordinary differential equations governing binomial moments, namely,
$\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{B}_{p,q}(t)$ $=\mathcal{B}_{p,q}(t)(D_{0}+D_{1}-up\bm{I}_{N}-\delta q\bm{I}_{N})+\mathcal{B}_{p-1,q}(t)D_{1}$ (21)
$+vp\mathcal{B}_{p,q-1}(t)+v(p+1)\mathcal{B}_{p+1,q-1}(t),\;\;p,q\in\mathbb{N}.$
Consider only the steady state, we may the time-derivatives in (21) to zero and thus obtain (8). Recall that $D_{0}+D_{1}=D$.
Appendix C Low-Order Binomial Moments
Based on the recurrence relation (8), analytical expressions for low-order binomial moments can be readily obtained, for example,
$\mathcal{B}_{1,0}=D_{1}(u\bm{I}_{N}-D)^{-1}$ (22)
$\mathcal{B}_{0,1}=vD_{1}(u\bm{I}_{N}-D)^{-1}(\delta\bm{I}_{N}-D)^{-1}$
$\mathcal{B}_{2,0}=D_{1}(u\bm{I}_{N}-D)^{-1}D_{1}(2u\bm{I}_{N}-D)^{-1}$
$\mathcal{B}_{1,1}=vD_{1}(u\bm{I}_{N}-D)^{-1}\left[I_{N}+(\delta\bm{I}_{N}-D)^{-1}D_{1}+2D_{1}(2u\bm{I}_{N}-D)^{-1}\right](u\bm{I}_{N}+\delta\bm{I}_{N}-D)^{-1}$
$\mathcal{B}_{0,2}=v^{2}D_{1}(u\bm{I}_{N}-D)^{-1}\left[I_{N}+(\delta\bm{I}_{N}-D)^{-1}D_{1}+2D_{1}(2u\bm{I}_{N}-D)^{-1}\right]$
$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\times(u\bm{I}_{N}+\delta\bm{I}_{N}-D)^{-1}(2\delta\bm{I}_{N}-D)^{-1}.$
To obtain the expressions in (10), recall that the $D$ is a $Q$-matrix, satisfying $D\bm{1}=\bm{0}_{N\times 1}$. Therefore,
$\left(k\bm{I}_{N}-D\right)^{-1}\bm{1}=\frac{1}{k}\bm{1},\;\;k>0.$ (23)
Using (23), it follows readily from (22) that
$B_{1,0}=\frac{1}{u}\bm{\pi}^{\top}D_{1}\bm{1},$ (24)
$B_{0,1}=\frac{v}{u\delta}\bm{\pi}^{\top}D_{1}\bm{1},$
$B_{2,0}=\frac{1}{2u}\bm{\pi}^{\top}D_{1}(u\bm{I}_{N}-D)^{-1}D_{1}\bm{1},$
$B_{1,1}=\frac{v}{u(u+\delta)}\bm{\pi}^{\top}D_{1}\bm{1}+\frac{v}{u+\delta}\bm{\pi}^{\top}D_{1}(u\bm{I}_{N}-D)^{-1}(\delta\bm{I}_{N}-D)^{-1}D_{1}\bm{1}$
$\;\;\;\;\;\;\;\;\;\;\;\;\;+\frac{v}{u(u+\delta)}\bm{\pi}^{\top}D_{1}(u\bm{I}_{N}-D)^{-1}D_{1}\bm{1},$
$B_{0,2}=\frac{v^{2}}{2u\delta(u+\delta)}\bm{\pi}^{\top}D_{1}\bm{1}+\frac{v^{2}}{2\delta(u+\delta)}\bm{\pi}^{\top}D_{1}(u\bm{I}_{N}-D)^{-1}(\delta\bm{I}_{N}-D)^{-1}D_{1}\bm{1}$
$\;\;\;\;\;\;\;\;\;\;\;\;\;+\frac{v^{2}}{2u\delta(u+\delta)}\bm{\pi}^{\top}D_{1}(u\bm{I}_{N}-D)^{-1}D_{1}\bm{1}.$
Appendix D Proof of (14)
$\lvert B_{0,2}-\widetilde{B_{2}}\rvert$ $=\lvert\frac{v^{2}}{2u\delta(u+\delta)}\bm{\pi}^{\top}D_{1}\bm{1}+\frac{v^{2}}{2\delta(u+\delta)}\bm{\pi}^{\top}D_{1}(u\bm{I}_{N}-D)^{-1}(\delta\bm{I}_{N}-D)^{-1}D_{1}\bm{1}$ (25)
$+\frac{v^{2}}{2u\delta(u+\delta)}\bm{\pi}^{\top}D_{1}(u\bm{I}_{N}-D)^{-1}D_{1}\bm{1}-\frac{v^{2}}{2u^{2}\delta}\bm{\pi}^{\top}D_{1}\bm{1}-\frac{v^{2}}{2u^{2}\delta}\bm{\pi}^{\top}D_{1}\left(\delta\bm{I}_{N}-D\right)^{-1}D_{1}\bm{1}\rvert$
$\leq\lvert\frac{v^{2}}{2u\delta(u+\delta)}\bm{\pi}^{\top}D_{1}\bm{1}-\frac{v^{2}}{2u^{2}\delta}\bm{\pi}^{\top}D_{1}\bm{1}\rvert$
$+\lvert\frac{v^{2}}{2\delta(u+\delta)}\bm{\pi}^{\top}D_{1}(u\bm{I}_{N}-D)^{-1}(\delta\bm{I}_{N}-D)^{-1}D_{1}\bm{1}-\frac{v^{2}}{2u^{2}\delta}\bm{\pi}^{\top}D_{1}\left(\delta\bm{I}_{N}-D\right)^{-1}D_{1}\bm{1}\rvert$
$+\frac{v^{2}}{2u\delta(u+\delta)}\lvert\bm{\pi}^{\top}D_{1}(u\bm{I}_{N}-D)^{-1}D_{1}\bm{1}\rvert$
$\leq\frac{v^{2}}{2u^{2}(u+\delta)}\lvert\bm{\pi}^{\top}D_{1}\bm{1}\rvert$
$+\frac{v^{2}}{2u(u+\delta)}\lvert\bm{\pi}^{\top}D_{1}(u\bm{I}_{N}-D)^{-1}(\delta\bm{I}_{N}-D)^{-1}D_{1}\bm{1}\rvert$
$+\frac{v^{2}}{2u\delta}\lvert\bm{\pi}^{\top}D_{1}(u\bm{I}_{N}-D)^{-1}(\delta\bm{I}_{N}-D)^{-1}D_{1}\bm{1}-\bm{\pi}^{\top}D_{1}\left(\frac{1}{u}\bm{I}_{N}\right)\left(\delta\bm{I}_{N}-D\right)^{-1}D_{1}\bm{1}\rvert$
$+\frac{v^{2}}{2u\delta(u+\delta)}\lvert\bm{\pi}^{\top}D_{1}(u\bm{I}_{N}-D)^{-1}D_{1}\bm{1}\rvert$
$\leq\frac{v^{2}}{2u^{2}(u+\delta)}\lVert\bm{\pi}\rVert_{1}\lVert D_{1}\rVert_{\infty}\lVert\bm{1}\rVert_{\infty}$
$+\frac{v^{2}}{2u(u+\delta)}\lVert\bm{\pi}\rVert_{1}\lVert D_{1}\rVert_{\infty}\lVert(u\bm{I}_{N}-D)^{-1}\rVert_{\infty}\lVert(\delta\bm{I}_{N}-D)^{-1}\rVert_{\infty}\lVert D_{1}\rVert_{\infty}\lVert\bm{1}\rVert_{\infty}$
$+\frac{v^{2}}{2u\delta}\lVert\bm{\pi}\rVert_{1}\lVert D_{1}\rVert_{\infty}\lVert(u\bm{I}_{N}-D)^{-1}-\frac{1}{u}\bm{I}_{N}\rVert_{\infty}\lVert(\delta\bm{I}_{N}-D)^{-1}\rVert_{\infty}\lVert D_{1}\rVert_{\infty}\lVert\bm{1}\rVert_{\infty}$
$+\frac{v^{2}}{2u\delta(u+\delta)}\lVert\bm{\pi}\rVert_{1}\lVert D_{1}\rVert_{\infty}\lVert(u\bm{I}_{N}-D)^{-1}\rVert_{\infty}\lVert D_{1}\rVert_{\infty}\lVert\bm{1}\rVert_{\infty}$
$\leq\frac{v^{2}}{2u^{2}(u+\delta)}\lVert D_{1}\rVert_{\infty}+\frac{v^{2}}{2u^{2}\delta(u+\delta)}\lVert D_{1}\rVert_{\infty}^{2}$
$+\frac{v^{2}}{2u^{2}\delta^{2}}\lVert D_{1}\rVert_{\infty}^{2}\lVert D\rVert_{\infty}+\frac{v^{2}}{2u^{2}\delta(u+\delta)}\lVert D_{1}\rVert_{\infty}^{2}$
$=\frac{v^{2}}{2u^{2}(u+\delta)}\lVert D_{1}\rVert_{\infty}+\frac{v^{2}}{u^{2}\delta(u+\delta)}\lVert D_{1}\rVert_{\infty}^{2}+\frac{v^{2}}{2u^{2}\delta^{2}}\lVert D_{1}\rVert_{\infty}^{2}\lVert D\rVert_{\infty}.$
In the penultimate inequality, we applied Hölder’s inequality. In the last inequality, we used an analogue to Theorem 4.1.2 in [MatrixComputations2013], stating that
THEOREM.
Let $C=(c_{i,j})_{N\times N}$ be a $N\times N$ strictly row diagonally dominant matrix with $\Theta:=\min_{1\leq i\leq N}\left(\mid c_{i,i}\mid-\sum_{j\neq i}\mid c_{i,j}\mid\right)$. Then $\lVert C^{-1}\rVert_{\infty}\leq\Theta^{-1}$.
Since $D$ is the generator of a continuous-time Markov chain, for any $\lambda>0$, the matrix $\lambda\bm{I}_{N}-D$ is strictly row diagonally dominant, with each row summing to $\lambda$. Therefore $\lVert\lambda\bm{I}_{N}-D\rVert_{\infty}\leq 1/\lambda$.
Note also that
$\lVert(u\bm{I}_{N}-D)^{-1}-\frac{1}{u}\bm{I}_{N}\rVert_{\infty}$ $=\lVert(u\bm{I}_{N}-D)^{-1}\left[\bm{I}_{N}-\frac{1}{u}(u\bm{I}_{N}-D)\right]\rVert_{\infty}$ (26)
$\leq\lVert(u\bm{I}_{N}-D)^{-1}\rVert_{\infty}\lVert D\rVert_{\infty}\leq\frac{1}{u}\lVert D\rVert_{\infty}.$
Appendix E Analytical Expression of Generating Function
In this section, we verify that (27) is the unique solution to (4) under initial condition $\mathcal{G}(z,w;0)=\bm{I}_{N}$, therefore presenting the explicit expression of generating function.
We claim that the unique solution to (4) under initial condition $\mathcal{G}(z,w;0)=\bm{I}_{N}$ is as follow.
$\mathcal{G}(z,w;t)=\bm{I}_{N}+\int_{0}^{t}H(t_{1})\mathrm{d}t_{1}+\sum_{k=2}^{\infty}\int_{0}^{t}\int_{0}^{t_{1}}\cdots\int_{0}^{t_{k-1}}H(t_{k})H(t_{k-1})\cdots H(t_{1})\mathrm{d}t_{k}\cdots\mathrm{d}t_{2}\mathrm{d}t_{1},$ (27)
where $H(s;z,w;t):=D_{0}+h(s;z,w,t)D_{1}\in\mathbb{R}^{N\times N}$ and $h(s;z,w;t)$ is a real-valued function whose expression is given below.
$h(s;z,w;t)$ $=\exp\!\left[us+\tfrac{v}{\delta}(1-w)\,\exp\!\bigl(\delta s-\delta t)\right]$ (28)
$\quad\times\Biggl\{z\exp\!\left[-ut-\tfrac{v}{\delta}(1-w)\right]$
$\qquad+\frac{u}{\delta}\left[\frac{v}{\delta}(1-w)\right]^{u/\delta}\exp(-ut)$
$\qquad\quad\times\Bigl[\Gamma\!\left(-\tfrac{u}{\delta};\tfrac{v}{\delta}(1-w)\exp(\delta s-\delta t)\right)-\Gamma\!\left(-\tfrac{u}{\delta};\tfrac{v}{\delta}(1-w)\right)\Bigr]\Biggr\}.$
In (28), $\Gamma(\gamma;x):=\int_{x}^{\infty}y^{\gamma-1}\mathrm{e}^{-y}\mathrm{d}y$ is the incomplete gamma function.
It is obvious that (27) satisfies the initial condition $\mathcal{G}(z,w;0)=\bm{I}_{N}$. Therefore we only need to verify that $\mathcal{G}(z,w;t)$ given by (27) satisfies the partial differential equation (4). The verification is straightforward but tedious, and we proceed in three steps.
E.1 Abbreviations
For simplicity, we first introduce some abbreviations for the calculation.
$a(z,w)=u(1-z)+vz(w-1),\;b(w)=\delta(1-w),$ (29)
$\theta=\text{any one of }\{z,w,t\},$
$c(w)=\frac{v}{\delta}(w-1),\;\alpha=-\frac{u}{\delta},\;X(s;w;t)=-c(w)e^{\delta(s-t)},$
$A(s;w;t)=\exp\big[us+X(s;w;t)\big],\;E(w;t)=\exp\big[-ut+c(w)\big],$
$M(w;t)=\frac{u}{\delta}\mathrm{e}^{-ut}[-c(w)]^{u/\delta},\;H(s;w;t)=\Gamma(\alpha;X(s;w;t))-\Gamma(\alpha,-c(w)).$
Note that with the above abbreviations, we have from (28)
$h(s;z,w;t)=A(s;w;t)\left[E(w;t)\,z+M(w;t)H(s;w;t)\right],0\leq s\leq t.$ (30)
E.2 Derivatives of Peano-Baker Series
We now take derivatives with respect to $t$ in (27), which takes the form of a Peano-Baker Series or time-ordering operator.
Note that
$\frac{\partial}{\partial t}\int_{0}^{t}H(t_{1})\mathrm{d}t_{1}=H(t)+\int_{0}^{t}\frac{\partial}{\partial t}H(t_{1})\mathrm{d}t_{1}=D_{0}+zD_{1}+\int_{0}^{t}\frac{\partial}{\partial t}H(t_{1})\mathrm{d}t_{1}.$ (31)
Similarly, for $k\geq 2$,
$\frac{\partial}{\partial t}\int_{0}^{t}\int_{0}^{t_{1}}\cdots\int_{0}^{t_{k-1}}H(t_{k})H(t_{k-1})\cdots H(t_{1})\mathrm{d}t_{k}\cdots\mathrm{d}t_{2}\mathrm{d}t_{1}$ (32)
$=\int_{0}^{t}\int_{0}^{t_{2}}\cdots\int_{0}^{t_{k-1}}H(t_{k})H(t_{k-1})\cdots H(t_{2})H(t)\mathrm{d}t_{k}\cdots\mathrm{d}t_{2}$
$+\int_{0}^{t}\frac{\partial}{\partial t}\left[\int_{0}^{t_{1}}\cdots\int_{0}^{t_{k-1}}H(t_{k})H(t_{k-1})\cdots H(t_{1})\mathrm{d}t_{k}\cdots\mathrm{d}t_{2}\right]\mathrm{d}t_{1}$
$=\int_{0}^{t}\int_{0}^{t_{2}}\cdots\int_{0}^{t_{k-1}}H(t_{k})H(t_{k-1})\cdots H(t_{2})\mathrm{d}t_{k}\cdots\mathrm{d}t_{2}(D_{0}+zD_{1})$
$+\int_{0}^{t}\int_{0}^{t_{1}}\cdots\int_{0}^{t_{k-1}}\frac{\partial}{\partial t}\left[H(t_{k})H(t_{k-1})\cdots H(t_{1})\right]\mathrm{d}t_{k}\cdots\mathrm{d}t_{2}\mathrm{d}t_{1}.$
In addition, for $k\geq 2$,
$\frac{\partial}{\partial\theta}\left[H(t_{k})H(t_{k-1})\cdots H(t_{1})\right]=$ $\left[\frac{\partial}{\partial\theta}H(t_{k})\right]H(t_{k-1})\cdots H(t_{1})+H(t_{k})\left[\frac{\partial}{\partial\theta}H(t_{k-1})\right]\cdots H(t_{1})$ (33)
$+\cdots+H(t_{k})H(t_{k-1})\cdots\left[\frac{\partial}{\partial\theta}H(t_{1})\right].$
Therefore,
$\frac{\partial}{\partial t}\mathcal{G}(z,w;t)$ $=\mathcal{G}(z,w;t)(D_{0}+zD_{1})$ (34)
$+\int_{0}^{t}\frac{\partial}{\partial t}H(t_{1})\mathrm{d}t_{1}+\sum_{k=2}^{\infty}\int_{0}^{t}\int_{0}^{t_{1}}\cdots\int_{0}^{t_{k-1}}\frac{\partial}{\partial t}\left[H(t_{k})H(t_{k-1})\cdots H(t_{1})\right]\mathrm{d}t_{k}\cdots\mathrm{d}t_{2}\mathrm{d}t_{1}.$
Using (33), the verification of the differential equation system (4) can be reduced to proving that the real-valued multi-variable function $h(s;z,w;t)$ satisfies the following differential equation.
$\partial_{t}h(s;z,w;t)=a(z,w)\partial_{z}h(s;z,w;t)+b(w)\partial_{w}h(s;z,w;t),0\leq s\leq t.$ (35)
In the following four subsections, we aim to prove (35) step by step.
E.3 Calculation of $\frac{\partial h(s;z,w;t)}{\partial t}$
In this subsection, we calculate $\frac{\partial h(s;z,w;t)}{\partial t}$ based on (30).
$\frac{\partial h(s;z,w;t)}{\partial t}=$ $\frac{\partial}{\partial t}A(s;w;t)\left[E(w;t)z+M(w;t)H(s;w;t)\right]$ (36)
$+A(s;w;t)\left[\frac{\partial}{\partial t}E(w;t)\,z+M(w;t)H(s;w;t)\right]$
$+A(s;w;t)\left[E(w;t)z+\frac{\partial}{\partial t}M(w;t)H(s;w;t)\right]$
$+A(s;w;t)\left[E(w;t)z+M(w;t)\frac{\partial}{\partial t}H(s;w;t)\right].$
Recall the definition of incomplete gamma function in $H(s;w;t)$, namely, $\Gamma(\alpha;x)=\int_{x}^{\infty}y^{\alpha-1}\mathrm{e}^{-y}\mathrm{d}y$. Through standard differentiation, we can readily obtain
$\frac{\partial}{\partial t}A(s;w;t)=A(s;w;t)c(w)\delta\mathrm{e}^{\delta(s-t)},$ (37)
$\frac{\partial}{\partial t}E(w;t)=-uE(w;t),$
$\frac{\partial}{\partial t}M(w;t)=-uM(w;t),$
$\frac{\partial}{\partial t}H(s;w;t)=-c(w)\delta\mathrm{e}^{\delta(s-t)}X(s;w;t)^{\alpha-1}\mathrm{e}^{-X(s;w;t)}.$
E.4 Calculation of $\frac{\partial h(s;z,w;t)}{\partial z}$
In this subsection, we calculate $\frac{\partial h(s;z,w;t)}{\partial z}$ based on (30).
$\frac{\partial h(s;z,w;t)}{\partial z}=$ $A(s;w;t)E(w;t).$ (38)
E.5 Calculation of $\frac{\partial h(s;z,w;t)}{\partial w}$
In this subsection, we calculate $\frac{\partial h(s;z,w;t)}{\partial w}$ based on (30).
$\frac{\partial h(s;z,w;t)}{\partial w}=$ $\frac{\partial}{\partial w}A(s;w;t)\left[E(w;t)z+M(w;t)H(s;w;t)\right]$ (39)
$+A(s;w;t)\left[\frac{\partial}{\partial w}E(w;t)\,z+M(w;t)H(s;w;t)\right]$
$+A(s;w;t)\left[E(w;t)z+\frac{\partial}{\partial w}M(w;t)H(s;w;t)\right]$
$+A(s;w;t)\left[E(w;t)z+M(w;t)\frac{\partial}{\partial w}H(s;w;t)\right].$
It can be easily calculated that
$\frac{\partial}{\partial w}A(s;w;t)=-A(s;w;t)\frac{v}{\delta}\mathrm{e}^{\delta(s-t)},$ (40)
$\frac{\partial}{\partial w}E(w;t)=\frac{v}{\delta}E(w;t),$
$\frac{\partial}{\partial w}M(w;t)=\frac{uv}{\delta^{2}c(w)}M(w;t),$
$\frac{\partial}{\partial w}H(s;w;t)=\frac{v}{\delta}\mathrm{e}^{\delta(s-t)}X(s;w;t)^{\alpha-1}\mathrm{e}^{-X(s;w;t)}-\frac{v}{\delta}[-c(w)]^{\alpha-1}\mathrm{e}^{c(w)}.$
E.6 The Partial Differential Equation Satisfied by $h(s;z,w;t)$
Assemble the results in (36), (38), and (39), one checks by direct (term-by-term) cancellation that (35) holds.
Appendix F Hierarchical solver for $\mathcal{B}_{p,q}$ by Layers
Algorithm 1 Hierarchical solver for $\mathcal{B}_{p,q}$ by Layers $L=p+q$
[…]
2:All $\{\mathcal{B}_{p,q}\}$ for layers $L=0,\dots,L_{\max}$.
3:function SolveRight(M,YM,\,Y) $\triangleright$ Return $X$ such that XM=YX\,M=Y (Use LU factorization [MatrixComputations2013])
4:  Factorize $M^{\top}=L\,U$
5:  Solve $L\,U\,X^{\top}=Y^{\top}$ for $X^{\top}$; return $X$
6:end function
[…]
8:for $L=1$ to $L_{\max}$ do $\triangleright$ Advance by layer
9:  // Previous-layer terms: $Y_{k}:=\mathcal{B}_{k,\,L-1-k}$ for $k=0,\dots,L-1$; set $Y_{-1}=Y_{L}=0$
10:  // Right-multiplication blocks: $M_{p}:=D-\big(up+\delta(L-p)\big)\bm{I}_{N}$ for $p=0,\dots,L$
11:  // Same-layer back-substitution (right $\to$ left): let $Z_{p}:=\mathcal{B}_{p,\,L-p}$
12:  $Z_{L}\leftarrow\textsc{SolveRight}(M_{n},\ -\,Y_{L-1}D_{1})$
13:  for p=L1p=L-1 downto 0 do
14:   $RHS\leftarrow-\,Y_{p-1}D_{1}\;-\;v\,p\,Y_{p}\;-\;v(p+1)\,Z_{p+1}$
15:   $Z_{p}\leftarrow\textsc{SolveRight}(M_{p},\ RHS)$
16:  end for
17:  // Write back this layer
18:  for $p=0$ to $L$ do
19:   $\mathcal{B}_{p,\,L-p}\leftarrow Z_{p}$
20:  end for
21:end for